Please help me with the 1H-NMR analysis of a molecule with the following data:

Formula: C14H10Cl4

δ 7.9 (doublet, 2H), 7.8 (doublet, 1H), 7.5 (doublet of doublets, 1H), 7.46 (doublet of doublets, 1H), 7.44 (doublet, 1H), 7.42 (doublet, 2H), 4.2 (doublet, 1H), 3.7 (doublet, 1H)

Please provide the fragment for each chemical shift
Answer #1

The structure; Gutio Cly DBE = Xeri (2+ Xc) (14H) (10 th) = 15- 7 - 8 The structure is thought to be, ya (Mito tare) a CO Ha